OpenStudy (migitmack):

How do I solve (2y+3)^2=9

OpenStudy (anonymous):

do you want a full answer...? or just a hint ?!

OpenStudy (migitmack):

A hiint would be nice

OpenStudy (migitmack):

It takes that long to type a hint?

OpenStudy (anonymous):

kk ! add -9 both sides you'll have \[(2y+3)^2-9=0\] and try to use this...\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (migitmack):

wher edid u get -9

OpenStudy (migitmack):



OpenStudy (anonymous):

kk \[(2y+1)^2-9=9+-9=0\] I just added -9 to each part of the equation !

OpenStudy (migitmack):